# Speculative Model Risk in Healthcare AI: Using Storytelling to Surface Unintended Harms

### v1

@@ Abstract
Artificial intelligence (AI) is rapidly transforming healthcare, enabling fast development of tools like stress monitors, wellness trackers, and mental health chatbots. However, rapid and low-barrier development can introduce risks of bias, privacy violations, and unequal access, especially when systems ignore real-world contexts and diverse user needs. Many recent methods use AI to detect risks automatically, but this can reduce human engagement in understanding how harms arise and who they affect. We present a human-centered framework that generates user stories and supports multi-agent discussions to help people think creatively about potential benefits and harms before deployment. In a user study, participants who read stories recognized a broader range of harms, distributing their responses more evenly across all 13 harm types. In contrast, those who did not read stories focused primarily on privacy and well-being (58.3%). Our findings show that storytelling helped participants speculate about a broader range of harms and benefits and think more creatively about AI’s impact on users. Dataset and code are available at https: https://anonymous.4open.science/r/storytelling-healthcare/README.md.
Speculative Model Risk in Healthcare AI: Using Storytelling to Surface Unintended Harms
Xingmeng Zhao, Dan Schumacher, Veronica Rammouz,  and Anthony Rios The University of Texas at San Antonio {xingmeng.zhao, anthony.rios}@utsa.edu
@@ 1 Introduction
AI is integrated into many aspects of daily life, including finance, healthcare, law, and recommendation systems Ashurst et al. (2020). In healthcare, tools such as stress monitors Kargarandehkordi et al. (2025), wellness trackers Fabbrizio et al. (2023), and mental health chatbots MacNeill et al. (2024) can directly influence well-being. Recent prompting methods like vibe coding Chow and Ng (2025) let people describe how an AI system should behave or sound, making it easier for non-experts to prototype AI tools. For instance, CareYaya allows caregivers to build stress-tracking applications using short natural language descriptions Kenny (2023). While this innovation speeds up AI development, it also introduces risks related to fairness, bias, and accountability Weidinger et al. (2023). These risks are serious in healthcare, where even small model errors can cause harm, including delayed treatment, privacy loss, or health inequities Roller et al. (2020); Chinta et al. (2025). Systems lacking safeguards for fairness and accountability may ultimately harm the very users they are meant to help Shelby et al. (2022, 2023). Governments are responding through measures such as the EU AI Act European Parliament (2023) and a U.S. Executive Order Biden (2023), which call for greater transparency, safety, and accountability Khan et al. (2025). Yet regulation remains region-specific and slow to adapt, making early ethical foresight essential for identifying risks, preventing misuse, and aligning AI systems with human values Saxena et al. (2025).
@@
Ethical challenges in AI are commonly addressed through two complementary strategies: documenting risks and anticipating them early. Model cards Mitchell et al. (2019) describe how a system works, its intended uses, and its limitations, and have been extended with interactive Crisan et al. (2022) and structured formats Bhat et al. (2023). RiskRAG Rao et al. (2025) builds on this idea by generating risk summaries from model cards and real-world incident data. While automation helps scale risk assessment, it can also reduce opportunities for thoughtful human reflection Kosmyna et al. (2025) and even introduce new harms Dutta et al. (2020). Another line of work focuses on anticipating potential misuse Herdel et al. (2024) and harm Deng et al. (2025); Saxena et al. (2025) early in the design process. Tools such as AHA! Buçinca et al. (2023), which uses storytelling, and Farsight Wang et al. (2024b), which provides real-time risk alerts, help surface ethical issues during ideation. Yet, as these systems increasingly automate ethical reflection, there is a risk that people may stop thinking critically about how AI systems could cause harm. When reflection is delegated to AI, users may begin to rely on its judgment instead of their own, making it harder to notice ethical or contextual issues.
@@
This issue is especially critical in healthcare, where small design mistakes can lead to serious harm Mennella et al. (2024); Gilbert et al. (2025). Many risks only appear after deployment, when AI systems face complex real-world conditions Mun et al. (2024); Kingsley et al. (2024). For example, mental health apps may fail to detect crises among certain groups, such as adolescents or non-native speakers Zhai et al. (2024). Current tools often rely on automation to predict such risks, but this can distance humans from ethical reasoning. Speculative design and design fiction offer a different approach, using imagined scenarios to explore how technologies might succeed or fail Rahwan et al. (2025). However, few approaches support human participation in ethical foresight or connect speculative thinking to real design workflows. Speculative storytelling bridges this gap by helping people reason about AI’s potential benefits and harms in context Li et al. (2025b).
@@
Building on Klassen and Fiesler (2022), who use speculative fiction to explore emerging technologies, we apply storytelling to prompt early ethical reflection in AI design (Figure 1). Our goal is to test whether storytelling helps people think more creatively about potential benefits and harms, encouraging human speculation rather than relying on AI to anticipate risks. We introduce a human-centered framework that combines automated user story generation with structured red-team discussions. Unlike plot-planning approaches Xie and Riedl (2024), our method produces context-sensitive stories grounded in users’ identities, behaviors, and needs. These stories help participants imagine how AI systems might succeed or fail in realistic scenarios, improving their ability to anticipate ethical and social implications. We use model cards to evaluate how clearly and diversely participants express potential harms. In our user study, story-driven discussions led to more context-specific risk identification and more detailed and varied responses to the model card.
@@
Our contributions are twofold. (1) We introduce a method that automatically generates user stories to help people imagine how an AI system could help or harm users before it is developed or deployed. (2) We present a user study showing story-driven discussions with AI agents help participants explore potential risks and benefits more creatively and think more broadly about ethical issues.
@@ 2 Related Work
Speculative Design. Speculative design uses imagined stories to explore how future technologies could affect people and society before those technologies exist. Rather than predicting the future, this approach uses what-if scenarios to prompt reflection on assumptions, values, and potential harms Klassen and Fiesler (2022); Hoang et al. (2018). This method treats fiction as a tool for early ethical reasoning, helping researchers consider social impacts while designing systems Rahwan et al. (2025). Recent work such as Buçinca et al. (2023) uses LLM-generated failure scenarios to automatically highlight harms, and Wang et al. (2024b) adds AI-driven risk prompts into prototyping workflows. In healthcare, Hoang et al. (2018) use “Fiction Probes” to let patients navigate imagined care scenarios, while Klassen and Fiesler (2022)’s Black Mirror Writers Room invites students to write near-future stories that reveal hidden harms. Beyond these examples, participatory workshops, crowdsourced case studies, and AI-assisted red-teaming extend speculative exploration by engaging a broader range of voices in imagining and testing possible futures Mun et al. (2024); Radharapu et al. (2023). Unlike prior work that uses AI to generate harms directly, we use AI-generated stories to spark human thinking and help people imagine harms themselves.
@@
Language-Based World Modeling. Humans often imagine situations to predict what may happen next, reason through alternatives, and make decisions Addis et al. (2009). Mental world modeling involves forming internal representations of objects, events, and relationships to simulate outcomes Johnson-Laird (1983). This process supports counterfactual and causal reasoning for planning and problem solving LeCun (2022). Large language models show related abilities through language. Xie et al. (2025) find that LLMs can act as world models that internally simulate state transitions, updating the situation step by step as events occur. Wang et al. (2023) show, through text games, that LLMs can translate user commands into coherent, evolving environments. Wang et al. (2024a) demonstrate that models can take on multiple reasoning personas that collaborate through internal dialogue to solve difficult tasks. Extending this idea, Fung et al. (2025) show that embodied agents use internal world models to predict surroundings, interpret user goals, and learn users’ mental models for better interaction. These studies present LLMs as text-based simulators that imagine, update, and reason about evolving environments. Building on this view, we frame story generation as language-based world imagination, where LLMs construct and simulate self-consistent worlds through narrative to reason about possible futures and their social, ethical, and technical implications.
@@
Figure 2: Overview of the Storytelling Framework. We first generate use case scenarios from AI concepts sourced from PubMed, Wired, and industry app descriptions. Next, we simulates role-playing and environment trajectories for each scenario, producing detailed simulation logs. Finally, we rephrase these logs into short stories that illustrate both potential benefits and harms of the AI system.
@@
Automated Story Generation. Early systems for automated story generation focused on modeling plot structure, often guided by narrative theories such as Propp’s functions to organize events Propp (1968); Yao et al. (2019). These approaches typically followed a two-step process: first, planning a sequence of key events, and then expanding them into detailed scenes Alhussain and Azmi (2021). With large language models, story generation has shifted toward unified systems where the model acts as both planner and writer. For example, Agents’ Room uses LLM-based character agents that jointly plan and enact stories Huot et al. , while Dramatron divides screenplay generation into loglines, character bios, plots, and dialogue Mirowski et al. (2023). HOLLMWOOD builds on this idea by using LLM role-play to create interactive, character-centered narratives Chen et al. (2024). Han et al. (2024) propose a Director–Actor framework for interactive scriptwriting, and Yu et al. (2025) extend this to a hierarchical approach that separates role-play and rewriting. To maintain both coherence and creativity, BookWorld introduces a world agent that tracks the story environment and coordinates character interactions within a global narrative plan Ran et al. (2025).
@@ 3 Methodology
Next, we used each specification as input to generate a set of ethically sensitive use-case scenarios. Each use case was represented as a 7-tuple $S=(a,u,s,x,b,h,f)$, where $a$ denoted the AI’s capability, $u$ the intended user (e.g., clinician), $s$ the subject (e.g., patient), $x$ the input or usage context, $b$ the expected benefit, $h$ the potential harm, and $f$ the failure trajectory (e.g., possible unintended or problematic uses) Shao et al. (2024). We used these structured representations to generate narrative user stories, which were then employed in red-teaming sessions to examine both user value and potential unintended harms. The full prompts used to extract model specifications and generate use cases are provided in Figure LABEL:sec:prompt-seed in Appendix.
@@ 4 Experiments
Dataset. We use GPT-4o to generate ethically sensitive use-case scenarios from 38 consumer health AI solutions sourced from Wired, industry product documentation, and PubMed. Each scenario acts as a narrative seed for simulation. For each AI concepts, we generate ten variations spanning different user roles (e.g., doctor, nurse, caregiver), settings (e.g., rural clinic, hospital, home), patient profiles (e.g., adolescent, older adult, multicultural family), and contextual conditions.
Baseline. We compare our method with a traditional plot-planning approach, where the model first outlines a plot before writing the story Yao et al. (2019); Xie and Riedl (2024). Using the same ethically sensitive seed, the baseline generates each story in a single step following a structured template. Each story consists of five sentences designed to prompt ethical reflection. The template directs the LLM to describe the AI system’s purpose, the people involved, the everyday use context, potential ethical risks, and how user identity may influence harm or misinterpretation. The full baseline prompt is provided in Figure 15 in Appendix.
Setting for Pairwise Comparison. We follow the evaluation setup from Li et al. (2025a) to assess story quality across multiple dimensions. Stories are evaluated according to five criteria: Creativity, measuring the originality and imagination of the plot and characters; Coherence, assessing narrative clarity and logical flow; Engagement, capturing how well the story maintains reader interest; Relevance, measuring consistency with the given prompt or scenario; and Likelihood of Harm or Benefit, evaluating whether the story depicts realistic AI behavior with meaningful social consequences. Following the arena-hard-auto evaluation method Li et al. , we use stories generated with the story-planning approach (by GPT-4o) as the reference baseline and compare them with stories produced by our method across different LLMs. For each metric, GPT-4o or human judges determine which story performs better or mark them as indistinguishable (“Tie”). Win rates are calculated based on these pairwise preferences, and the full configuration details and evaluation prompts are included in the Appendix. To eliminate positional bias, we randomize the order of story pairs and alternate their positions across comparisons. See Figures 16 and 17 for detailed criteria in Appendix.
@@
LLM-as-a-Judge Evaluation. As shown in Table 3, our Storytelling method consistently outperforms all baselines across every metric. When combined with the Gemma model, it achieves win rates of 89.45% for creativity, 92.15% for coherence, 92.75% for engagement, 85.65% for relevance, and 96.05% for likelihood, yielding an overall average of 91.21%. In contrast, the baseline Gemma records 72.76%, and Llama3 reaches 69.71%, indicating gains of roughly +15–25 points across dimensions. We use GPT-4o as the evaluator with the temperature set to 0.1 to ensure deterministic and consistent judgments across comparisons. Comparing models, baseline Gemma slightly outperforms Llama3 in most metrics, but under the Storytelling framework, Llama3 nearly closes the gap with an overall score of 89.24%. Interestingly, Llama3 surpasses Gemma in coherence (94.75 vs. 92.15) and performs equally well in relevance (85.65), suggesting that Llama3 shows stronger structural reasoning and coherence, while Gemma excels in narrative creativity and expressiveness. Overall, these results demonstrate that integrating world and role-based modeling enables models to reason about events, sustain coherent narratives, and produce stories that are both imaginative and believable.
@@
Human Evaluation. To complement the LLM-as-a-Judge evaluations and reduce potential bias, we conducted human preference evaluations. Two graduate student annotators independently evaluated 100 story pairs for each model and method. As shown in Figure 3, our Storytelling method is consistently preferred over all baselines, achieving 88% preference for Llama3 and 76% for Gemma, patterns that align with GPT-4o based evaluations. Notably, human judges show a slightly stronger preference for Llama3, suggesting it produces stories that are easier to follow and more engaging, while Gemma tends to generate more expressive and stylistically rich narratives. We further measure inter-annotator consistency using Cohen’s kappa Cohen (1960). As shown in Table 4, agreement scores range from 0.619 to 0.729 across models and methods, indicating substantial reliability. Even with small differences, LLM-as-a-Judge and human evaluations show that Storytelling performs well in both settings.

### v3

# Telling Speculative Stories to Help Humans Imagine the Harms of Healthcare AI

@@ Abstract
Artificial intelligence (AI) is rapidly transforming healthcare, enabling fast development of tools like stress monitors, wellness trackers, and mental health chatbots. However, rapid and low-barrier development can introduce risks of bias, privacy violations, and unequal access, especially when systems ignore real-world contexts and diverse user needs. Many recent methods use AI to detect risks automatically, but this can reduce human engagement in understanding how harms arise and who they affect. We present a human-centered framework that generates user stories and supports multi-agent discussions to help people think creatively about potential benefits and harms before deployment. In a user study, participants who read stories recognized a broader range of harms, distributing their responses more evenly across all 17 harm types. In contrast, those who did not read stories focused primarily on privacy and well-being (79.1%). Our findings show that storytelling helped participants speculate about a broader range of harms and benefits and think more creatively about AI’s impact on users. Dataset and code are available at https: https://anonymous.4open.science/r/storytelling-healthcare/README.md.
Telling Speculative Stories to Help Humans Imagine the Harms of Healthcare AI
Xingmeng Zhao, Tongnian Wang, Dan Schumacher, Veronica Rammouz,  and Anthony Rios The University of Texas at San Antonio {xingmeng.zhao, anthony.rios}@utsa.edu
@@ 1 Introduction
Artificial intelligence (AI) is increasingly embedded in everyday domains such as finance, healthcare, and law Ashurst et al. (2020). In healthcare, AI tools including stress monitors Kargarandehkordi et al. (2025), wellness trackers Fabbrizio et al. (2023), and mental health chatbots MacNeill et al. (2024) can directly affect users’ well-being. New prompting approaches such as vibe codingChow and Ng (2025) allow non-experts to describe desired system behavior in natural language, enabling rapid prototyping of AI applications, for example through platforms like CareYayaKenny (2023). However, these developments introduce risks related to fairness, bias, and accountability Weidinger et al. (2023), which are especially critical in healthcare, where small errors can cause harm, including delayed treatment, privacy loss, and health inequities Roller et al. (2020); Chinta et al. (2025). AI systems without appropriate safeguards may ultimately harm the users they aim to support Shelby et al. (2022, 2023). Although governments have begun responding through efforts such as the EU AI Act European Parliament (2023) and a U.S. Executive Order Biden (2023), which emphasize transparency and accountability Khan et al. (2025), regulation remains slow and fragmented, making early ethical foresight essential for aligning AI systems with human values Saxena et al. (2025).
@@
Ethical challenges in AI are commonly addressed through two complementary approaches: documenting known risks and anticipating potential harms early in design. Model cards Mitchell et al. (2019) describe a system’s purpose, behavior, and limitations, and have been extended with interactive Crisan et al. (2022) and structured formats Bhat et al. (2023). Building on this work, RiskRAG Rao et al. (2025) automatically generates risk summaries from model cards and real-world incident data. While such automation helps scale ethical assessment, it may reduce opportunities for human reflection Kosmyna et al. (2025) and introduce new harms Dutta et al. (2020). Another line of research focuses on anticipating misuse Herdel et al. (2024) and harm Deng et al. (2025); Saxena et al. (2025) early in the design process, using tools such as AHA!Buçinca et al. (2023) and FarsightWang et al. (2024b). However, as ethical reflection becomes increasingly automated, users may rely on AI judgment rather than their own, making ethical and contextual issues harder to recognize.
@@
This challenge is particularly critical in healthcare, where small design mistakes can cause serious harm Mennella et al. (2024); Gilbert et al. (2025). Many risks only emerge after deployment, when AI systems operate in complex real-world settings Mun et al. (2024); Kingsley et al. (2024). For example, mental health applications may fail to detect crises for certain populations, such as adolescents or non-native speakers Zhai et al. (2024). Although existing tools often rely on automation to predict these risks, this can distance humans from ethical reasoning. Speculative design and design fiction offer an alternative by using imagined scenarios to explore how technologies might succeed or fail Rahwan et al. (2025). However, few approaches meaningfully support human participation in ethical foresight or integrate speculative thinking into real design workflows. Speculative storytelling addresses this gap by helping people reason about AI’s potential benefits and harms within realistic contexts Li et al. (2025b).
@@
Building on Klassen and Fiesler (2022), who use speculative fiction to examine emerging technologies, we apply storytelling to prompt early ethical reflection in AI design (Figure 1). Our approach tests whether stories encourage creative human speculation about potential benefits and harms, rather than relying on AI to anticipate risks. We introduce a human-centered framework that combines automated user story generation with structured red-team discussions. Unlike plot-planning methods Xie and Riedl (2024), our approach generates context-sensitive stories grounded in users’ identities, behaviors, and needs. These stories help participants envision realistic success and failure scenarios, improving their ability to identify ethical and social risks. Using model cards as an evaluation tool, we show that story-driven discussions lead to more context-specific, detailed, and diverse expressions of potential harms.
@@
Our contributions are twofold. (1) We introduce a human-centered method that automatically generates context-sensitive user stories to help people imagine how an AI system could help or harm users before it is developed or deployed. (2) We present a user study showing story-driven discussions with AI agents help participants explore potential risks and benefits more creatively and think more broadly about ethical issues.
@@ 2 Related Work
Speculative Design. Speculative design uses imagined scenarios to explore how future technologies might affect people and society before they are built. Rather than predicting outcomes, it relies on what-if stories to prompt reflection on assumptions, values, and potential harms Klassen and Fiesler (2022); Hoang et al. (2018), treating fiction as a tool for early ethical reasoning Rahwan et al. (2025). Prior work has applied this approach using AI-generated failure cases Buçinca et al. (2023), risk prompts embedded in prototyping workflows Wang et al. (2024b), and participatory methods, like Fiction Probes in healthcare Hoang et al. (2018) and the Black Mirror Writers Room Klassen and Fiesler (2022). Other extensions include participatory workshops, crowdsourced case studies, and AI-assisted red-teaming to broaden speculative exploration Mun et al. (2024); Radharapu et al. (2023), as well as non-narrative artifacts like generated comments or judgments Ballard et al. (2019). In contrast to directly generating harms, our approach uses AI-generated stories to prompt human reflection and help participants imagine harms.
@@
Language-Based World Modeling. Humans imagine situations to anticipate outcomes, explore alternatives, and guide decisions Addis et al. (2009). This ability relies on mental world modeling, in which people form internal representations of objects, events, and relationships to simulate possible futures Johnson-Laird (1983), supporting causal and counterfactual reasoning for planning and problem solving LeCun (2022). Recent work shows large language models (LLMs) exhibit related capabilities through language. LLMs can act as text-based world models that simulate state transitions over time Xie et al. (2025), generate coherent and evolving environments in response to user actions Wang et al. (2023), and reason through internal multi-persona dialogue Wang et al. (2024a). Embodied agents further extend this idea by using internal world models to predict environments, infer user goals, and adapt to users’ mental models Fung et al. (2025). Building on this perspective, we frame story generation as language-based world imagination, where LLMs construct self-consistent narrative worlds to reason about possible futures and their social, ethical, and technical implications.
@@
Figure 2: Overview of the Storytelling Framework. We first generate use case scenarios from AI concepts sourced from PubMed, Wired, and industry app descriptions. Next, we simulate role-playing and environment trajectories for each scenario, producing detailed simulation logs. Finally, we rephrase these logs into short stories that illustrate both potential benefits and harms of the AI system.
@@
Automated Story Generation. Early work on automated story generation emphasized explicit plot modeling, often drawing on narrative theories such as Propp’s functions to structure events, and typically adopted a two-stage pipeline that first planned key events and then realized them as full scenes Propp (1968); Alhussain and Azmi (2021). With LLMs, this paradigm has shifted toward unified frameworks in which a single model jointly plans and writes narratives: Agents’ Room coordinates LLM-based character agents for collaborative story enactment Huot et al. , Dramatron decomposes screenplay generation into structured components such as loglines and dialogue Mirowski et al. (2023), and HOLLMWOOD uses LLM role-play to produce interactive, character-centered stories Chen et al. (2024). Han et al. (2024) introduce a Director–Actor framework for interactive scriptwriting, which Yu et al. (2025) extend with hierarchical role separation, while BookWorld adds a world agent to track global state and balance coherence with creativity Ran et al. (2025).
@@ 3 Methodology
Next, we used each specification as input to generate a set of ethically sensitive use-case scenarios. Each use case was represented as a 7-tuple $S=(a,u,s,x,b,h,f)$, where $a$ denoted the AI’s capability, $u$ the intended user (e.g., clinician), $s$ the subject (e.g., patient), $x$ the input or usage context, $b$ the expected benefit, $h$ the potential harm, and $f$ the failure trajectory (e.g., possible unintended or problematic uses) Shao et al. (2024). We used these structured representations to generate narrative user stories, which were then employed in red-teaming sessions to examine both user value and potential unintended harms. The full prompts used to extract model specifications and generate use cases are provided in Figure 12 in Appendix.
@@ 4 Experiments
Dataset. We used GPT-4o to generate ethically sensitive use-case scenarios from 38 consumer health AI solutions sourced from Wired, industry product documentation, and PubMed. Each scenario acted as a narrative seed for simulation. For each AI concepts, we generated ten variations spanning different user roles (e.g., doctor, nurse, caregiver), settings (e.g., rural clinic, hospital, home), patient profiles (e.g., adolescent, older adult, multicultural family), and contextual conditions.
Baseline. We compared our method with a traditional plot-planning approach, where the model first outlines a plot before writing the story Yao et al. (2019); Xie and Riedl (2024). Using the same ethically sensitive seed, the baseline generated each story in a single step following a structured template. Each story consisted of five sentences designed to prompt ethical reflection. The template directs the LLM to describe the AI system’s purpose, the people involved, the everyday use context, potential ethical risks, and how user identity may influence harm or misinterpretation. The full baseline prompt is provided in Figure 15 in Appendix.
Setting for Pairwise Comparison. We followed the evaluation setup from Li et al. (2025a) to assess story quality across multiple dimensions. Stories were evaluated according to five criteria: Creativity, measuring the originality and imagination of the plot and characters; Coherence, assessing narrative clarity and logical flow; Engagement, capturing how well the story maintains reader interest; Relevance, measuring consistency with the given prompt or scenario; and Likelihood of Harm or Benefit, evaluating whether the story depicts realistic AI behavior with meaningful social consequences. Following the arena-hard-auto evaluation method Li et al. , we used stories generated with the story-planning approach (by GPT-4o) as the reference baseline and compared them with stories produced by our method across different LLMs. For each metric, GPT-4o or human judges determined which story performs better or mark them as indistinguishable (“Tie”). Win rates were calculated based on these pairwise preferences, and the full configuration details and evaluation prompts are included in the Appendix. To eliminate positional bias, we randomized the order of story pairs and alternate their positions across comparisons. See Figures 16 and 17 for detailed criteria in Appendix.
@@
LLM-as-a-Judge Evaluation. As shown in Table 3, our Storytelling method consistently outperforms all baselines across every metric. When combined with the Gemma model, it achieves win rates of 89.45% for creativity, 92.15% for coherence, 92.75% for engagement, 85.65% for relevance, and 96.05% for likelihood, yielding an overall average of 91.21%. In contrast, the baseline Gemma records 72.76%, and Llama3 reaches 69.71%, indicating gains of roughly +15–25 points across dimensions. We use GPT-4o as the evaluator with the temperature set to 0.1 to ensure deterministic and consistent judgments across comparisons. Comparing models, baseline Gemma slightly outperforms Llama3 in most metrics, but under the Storytelling framework, Llama3 nearly closes the gap with an overall score of 89.24%. Interestingly, Llama3 surpasses Gemma in coherence (94.75 vs. 92.15) and performs equally well in relevance (85.65), suggesting that Llama3 shows stronger structural reasoning and coherence, while Gemma excels in narrative creativity and expressiveness. Overall, these results demonstrate that integrating world and role-based modeling enables models to reason about events, sustain coherent narratives, and produce stories that are both imaginative and believable. For robustness, we report results from two additional LLM-as-a-judge models in the appendix A.8.
@@
Human Evaluation. To complement the LLM-as-a-Judge evaluations and reduce potential bias, we conducted human preference evaluations. Two graduate student annotators independently evaluated 100 story pairs for each model and method. As shown in Figure 3, our Storytelling method is consistently preferred over all baselines, achieving 88% preference for Llama3 and 76% for Gemma, patterns that align with GPT-4o based evaluations. Notably, human judges show a slightly stronger preference for Llama3, suggesting it produces stories that are easier to follow and more engaging, while Gemma tends to generate more expressive and stylistically rich narratives. We further measure inter-annotator consistency using Cohen’s kappa Cohen (1960). As shown in Table 4, agreement scores range from 0.619 to 0.729 across models and methods, indicating substantial reliability. Overall, both human and LLM evaluations consistently agree that Storytelling outperforms all baselines, and differences between Gemma and Llama3 reflect LLM preference for detail versus human preference for clarity.